# Diff3DS: Generating View-Consistent 3D Sketch via Differentiable Curve Rendering

### v1

@@ Abstract
3D sketches are widely used for visually representing the 3D shape and structure of objects or scenes. However, the creation of 3D sketch often requires users to possess professional artistic skills. Existing research efforts primarily focus on enhancing the ability of interactive sketch generation in 3D virtual systems. In this work, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. Specifically, we perform perspective projection to render the 3D rational Bézier curves into 2D curves, which are subsequently converted to a 2D raster image via our customized differentiable rasterizer. Our framework bridges the domains of 3D sketch and raster image, achieving end-to-end optimization of 3D sketch through gradients computed in the 2D image domain. Our Diff3DS can enable a series of novel 3D sketch generation tasks, including text-to-3D sketch and image-to-3D sketch, supported by the popular distillation-based supervision, such as Score Distillation Sampling (SDS). Extensive experiments have yielded promising results and demonstrated the potential of our framework. Project: https://yiboz2001.github.io/Diff3DS/
@@ 1 Introduction
3D sketches, which utilize strokes to emphasize abstraction and visually encapsulate the 3D shape and structure of objects or scenes, serve as indispensable tools for visualizing concepts and ideas. Existing research efforts have primarily focused on exploring how to enable interactive creation of 3D sketch in 3D virtual systems [44, 12, 5] and immersive environments [53, 48, 19, 17, 18, 54, 4, 55]. Such interactive 3D sketch creation tools provide artists with unprecedented freedom, allowing them to draw their inspirations directly in the immersive 3D environment. However, it usually requires users to possess professional artistic skills and experience for 3D operation, making the 3D sketch creation less friendly for ordinary users. There is a notable absence of a user-friendly 3D sketch generation method in the community.
@@
In recent years, content reconstruction or generation via differentiable rendering has gained great attention in the fields of computer graphics and vision research. The goal is to convert the traditional rendering pipeline into a differentiable image synthesis process and employ the 2D image supervision to optimize the object or scene representation. For example, NeRF [24] bridges the 3D neural representation and raster image domain with a differentiable volume rendering pipeline, and has contributed to the further development of the 3D generation [25, 38, 15, 28, 21] by leveraging the continuous advancements in multimodal supervision [32, 28]. For the generation of vectorized or parametric content, differentiable rendering has also been investigated [20, 33, 10, 16, 49, 31, 50, 56, 6, 40, 39]. One representative work is DiffVG [20], which proposes a differentiable 2D vector graphics rasterizer to compute gradients from raster images and optimize the vectorized image (e.g., 2D Bézier curves and other parametric shapes) via back-propagation. For generation of sketch, CLIPasso generates 2D sketches using a differentiable rasterizer to optimize the parameters of the Bézier curves directly with respect to a CLIP-based perceptual loss. Recently, 3Doodle [8] generates view-consistent 3D sketches of the target object by directly optimizing the parameters of 3D strokes to minimize the perceptual losses for given multi-view images.
@@
Although 3Doodle achieves promising results for generating 3D sketches, it has two main limitations. First, 3Doodle performs orthogonal projection to obtain the 2D sketch rendering and ignores the depth order of the curves, which may lead to errors in the 3D sketch. Second, its optimization requires supervision from multiview images, which makes it more like a reconstruction pipeline instead of a more flexible generation pipeline. In this paper, we aim to solving above issues and present Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. Specifically, we represent the 3D sketch as a set of 3D rational Bézier curves and perform the perspective projection to obtain the 2D rational Bézier curves. Then, we propose a new differentiable rasterizer based on DiffVG [20] to accurately render the projected 2D curves based on the depth order. Our framework supports end-to-end optimization of 3D curve primitives under flexible supervisions such as the distillation-based loss. By employing the recent Score Distillation Sampling (SDS) [28, 22] to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the text or single image input. We conduct quantitative and qualitative comparisons between Diff3DS and related methods, and the results show the superiority of our approach in generating view-consistent 3D sketch. Ablation studies also verify the effectiveness our our key hyperparameters and components.
@@
We summarize our contributions as follows: (1) We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. (2) We are the first to represent the 3D curve as 3D rational Bézier curve and design a new differentiable rasterizer to render the perspectively projected 2D rational Bézier curves, while considering the depth order. (3) We conduct comprehensive experiments on novel text-to-3D and image-to-3D sketch generation tasks and the results demonstrate the superiority of our method.
@@ 2 Related Works
Existing research on 3D sketch generation mainly focus on interactive artistic sketch creation within 3D virtual systems [44, 12, 5], and immersive environments [53, 48, 19, 17, 18, 54, 4, 55]. However, these works usually require users to possess professional artistic skills and 3D operation experience, making it less user-friendly for ordinary users. The recent 3Doodle [8] is able to generate expressive 3D sketches from multiview image observations, but it is more like multiview reconstruction instead of generation and the quality of the generated results highly depends on the number of observation viewpoints. By distilling prior knowledge from pre-trained 2D image generation models using methods like Score Distillation Sampling [28, 22], we achieve user-friendly and view-consistent 3D sketch generation from flexible text or single image input.
@@
There have been several works for differentiable rendering of curves. DiffVG [20] proposes a differentiable rasterizer for the creation of vector graphics represented by 2D curves and shapes. DRPG [46] investigates differentiable rendering of 3D parametric curves and surfaces, by piecewisely approximating the continuous parametric representations with a triangle mesh. The 3Doodle [8] shares a similar differentiable rendering pipeline with us, including projecting the 3D Bézier curves to 2D Bézier curves, which are then rendered by the differentiable rasterizer from DiffVG. However, 3Doodle performs orthogonal projection to obtain 2D Bézier curves and ignores the depth ordering between the curves. Instead, we perform perspective projection to the 3D rational Bézier curves and designs a new DiffVG-based ratersizer which can maintain the depth ordering between the projected 2D rational Bézier curves.
@@
Motivated by the success of the text-to-image research, pioneering works [28, 41] introduce the Score Distillation Sampling (SDS) algorithm and leverage the pre-trained text-to-image models as prior knowledge to optimize 3D representations. Subsequent works combine SDS with various differentiable 3D representations to explore the capabilities of text-to-3D object generation, such as DMTet [21, 7] and 3D Gaussians [37, 52]. Furthermore, Zero-1-to-3 [22] proposes an image-guided SDS algorithm that distills the 3D-consistent priors to optimize 3D representations given the input image, and its pipeline inspires the development of subsequent image-to-3D object methods [22, 30, 42, 47]. Distinguishing from prior endeavors, our work stands out by focusing on generating view-consistent 3D sketch by optimizing 3D parametric curve primitives.
@@ 3 Preliminary
DiffVG [20] proposes a differentiable rasterizer for vector graphics that supports converting 2D Bézier curves to the raster image domain and back-propagating the gradient for optimization.
Given the 2D curves parameters $\Theta$, the vector graphics scene is defined as $f(x,y,\Theta)$, and the raster image is defined as the 2D grid sampling over the space of $f(x,y,\Theta)$. To compute the color of a 2D location $(x,y)\in\mathbb{R}^{2}$, DiffVG utilizes the inside-outside test [26] to find the curves overlapped with the location initially. Subsequently, it sorts them according to a user-specified order and calculates the color using alpha blending [29]. Due to the inside-outside test, the scene function $f$ is not differentiable with respect to curve parameters. Thus, it uses the anti-aliasing technique to make the pixel color differentiable. By prefiltering $f$ over a convolution kernel $k$ with support $A$, sampling $I(x,y)$ at pixels can yield an alias-free image:
@@
DiffVG proposes two approaches Monte Carlo Sampling and Analytical Prefiltering to evaluate the pixel integral, which does not have a closed-form solution in general. We focus on the first approach which performs better. The pixel integral can be discretized with Monte Carlo sampling [11]:
@@ 4.1 Overview
We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves. A 3D sketch $\tilde{\Theta}$ is defined as a collection of 3D parametric curve strokes that reside in 3D space. We particularly focus on the 3D rational Bézier curve and our derivation supports the 3D linear, quadratic, and cubic rational Bézier curves. The theoretical trainable parameters include the control point position, curve width, and rgb-alpha color. Follow the previous methods [40, 39, 8], we represent each stroke by a 3D cubic Bézier curve and only optimize the control point position in our current implementation. Moreover, weights for rational Bézier curves are disregarded to simplify the computation. Fig. 2 illustrates the pipeline.
@@
In this section, we mainly explain how to perspectively project the 3D rational Bézier curves in camera coordinates onto the 2D camera focal plane. Given the control points $\tilde{P}_{i}=(\tilde{P}_{i}^{x},\tilde{P}_{i}^{y},\tilde{P}_{i}^{z})\in\mathbb{R}^{3}$, the associated non-negative weights $\tilde{w}_{i}$ and the Bernstein polynomials $B_{i,n}(t)$, for $0\leq t\leq 1$, define the 3D rational Bézier curve $\tilde{p}(t)$ of degree $n$ by:
@@
Due to rational Bézier curves are projective invariant, the curve $p(t)$ still is a 2D rational Bézier curve defined by projected control points. Given projected 2D control points $P_{i}=(P_{i}^{x},P_{i}^{y})=(\frac{\tilde{P}_{i}^{x}}{\tilde{P}_{i}^{z}}f,\frac{\tilde{P}_{i}^{y}}{\tilde{P}_{i}^{z}}f)\in\mathbb{R}^{2}$ and adjusted weights $w_{i}=\tilde{w}_{i}\tilde{P}_{i}^{z}$, $p(t)$ can be written as following:
@@
We provide the detailed proof in Appendix B.
Figure 2: We generate the 3D sketch $\tilde{\Theta}$, represented as a set of 3D strokes, from the text or image input. We render the raster image $I$ from the random camera via our differentiable renderer (Sec. 4). Then, the pre-trained diffusion model, conditioned on the input, diffuses the rendering $I$ and predicts the pseudo ground truth $\hat{I}_{0}$. The discrepancies between $\hat{I}_{0}$ and $I$ are used to update the 3D sketch.
@@
This section primarily focuses on the differentiable rasterization of 2D rational Bézier curves $\Theta$ obtained through projection.
To render the projected curves and maintain their depth orders, we customize a new differentiable rasterizer based on DiffVG [20]. Given a 2D pixel location $(x,y)$ and projected curves $\Theta$, we employ the inside-outside test to identify all the overlapping curves with the location. The inside-outside test identifies a curve as overlapping with a location if the closest distance between them is less than half of the curve width, and the closest point on each overlapping curve with the location are considered as the overlapping point. Then, we sort the overlapping point to calculate the color of the pixel location $(x,y)$ using alpha blending [29]. Our sort order follows the depth order between curve primitives. We employ the unprojected z-depth in the camera space of the overlapping point as the primary sort order. To mitigate the z-fighting issue [^1] https://en.wikipedia.org/wiki/Z-fighting caused by floating-point error, we also employ a user-specified order as the secondary order.
Identifying the overlapping curve is challenging. For the $n$th-degree rational curve $p(t)$, computing the $t$ that minimizes the distance $(p(t)-q)^{2}$ where $q=(x,y)$ equals to solve the roots of a polynomial with degree $23n-$ 2. For the cubic rational curve, a 7th order polynomial needs to be solved, which doesn’t have a closed-form solution. Inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method [45]. The iterative solver obtains its initial guess from isolator polynomials [34] – the real roots of a 7th order polynomial can be isolated by the roots of two cubic auxiliary polynomials. Please refer to Appendix C for more details.
After obtaining the 2D curve scene function $f(x,y,\Theta)$, the rendered raster image $I$ will be calculated using the Monte Carlo Sampling strategy (Sec. 3) with Eq. 3. Fig. 3 shows an example.
Figure 3: A rasterized result contains 3 quadratic rational Bézier curves and a line. All curves share the same control point positions in pixel space and the depths but with different weights. Our rasterizer accurately renders the curves and faithfully maintains occlusions according to the depth order. The user-specified sort order results in the pixel at position (1023, 0) being the color of Curve C, as it has a higher priority than the other overlapping curve points with the same depth.

### v2

@@ Abstract
3D sketches are widely used for visually representing the 3D shape and structure of objects or scenes. However, the creation of 3D sketch often requires users to possess professional artistic skills. Existing research efforts primarily focus on enhancing the ability of interactive sketch generation in 3D virtual systems. In this work, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. Specifically, we perform perspective projection to render the 3D rational Bézier curves into 2D curves, which are subsequently converted to a 2D raster image via our customized differentiable rasterizer. Our framework bridges the domains of 3D sketch and raster image, achieving end-to-end optimization of 3D sketch through gradients computed in the 2D image domain. Diff3DS can enable a series of novel 3D sketch generation tasks, including text-to-3D sketch and image-to-3D sketch, supported by the popular distillation-based supervision, such as Score Distillation Sampling. Extensive experiments have yielded promising results and demonstrated the potential of our framework. Project page is at https://yiboz2001.github.io/Diff3DS/
@@ 1 Introduction
3D sketches, which utilize strokes to emphasize abstraction and visually encapsulate the 3D shape and structure of objects or scenes, serve as indispensable tools for visualizing concepts and ideas. Existing research efforts have primarily focused on exploring how to enable interactive creation of 3D sketch in 3D virtual systems (Wesche & Seidel, 2001; Grimm & Joshi, 2012; Bae et al., 2008) and immersive environments (Yu et al., 2021a; Xin et al., 2008; Kim et al., 2023; Jiang et al., 2021; Kim & Bae, 2016; Yu et al., 2021b; Arora & Singh, 2021; Yue et al., 2017). Such interactive 3D sketch creation tools provide artists with unprecedented freedom, allowing them to draw their inspirations directly in the immersive 3D environment. However, it usually requires users to possess professional artistic skills and experience for 3D operation, making the 3D sketch creation less friendly for ordinary users. There is a notable absence of a user-friendly 3D sketch generation method in the community. Moreover, in the area of wire art generation, 3D sketch or 3D curves are also widely studied to abstract the desired visual concepts from diverse inputs, e.g., 3D surfaces (Yang et al., 2021), multi-view images (Liu et al., 2017) and text (Tojo et al., 2024; Qu et al., 2024). These representations provide essential shape and structure of the artwork as well as hold significant potential as intermediate formats for tasks such as 3D reconstruction. However, generating view-consistent 3D sketches from flexible inputs, such as text or a single image, remains largely unexplored.
@@
In recent years, content reconstruction or generation via differentiable rendering has gained great attention in the fields of computer graphics and vision research. The goal is to convert the traditional rendering pipeline into a differentiable image synthesis process and employ the 2D image supervision to optimize the object or scene representation. For example, NeRF (Mildenhall et al., 2021) bridges the 3D neural representation and raster image domain with a differentiable volume rendering pipeline, and has contributed to the further development of the 3D generation (Mohammad Khalid et al., 2022; Tsalicoglou et al., 2024; Jain et al., 2022; Poole et al., 2023; Lin et al., 2023) by leveraging the continuous advancements in multimodal supervision (Radford et al., 2021; Poole et al., 2023). For the generation of vectorized or parametric content, differentiable rendering has also been investigated (Li et al., 2020; Schaldenbrand et al., 2022; Frans et al., 2022; Jain et al., 2023; Xing et al., 2023; Qu et al., 2023; Xing et al., 2024; Banerjee et al., 2024; Vinker et al., 2022; 2023; Thamizharasan et al., 2024b; a). One representative work is DiffVG (Li et al., 2020), which proposes a differentiable 2D vector graphics rasterizer to compute gradients from raster images and optimize the vectorized image (e.g., 2D Bézier curves and other parametric shapes) via back-propagation. For generation of sketch, CLIPasso generates 2D sketches using DiffVG to optimize the parameters of the 2D Bézier curves directly with respect to a CLIP-based perceptual loss. Recently, 3Doodle (Choi et al., 2024) generates view-consistent 3D sketches of the target object by directly optimizing the parameters of 3D strokes to minimize the perceptual losses for given multi-view images. It also proposes a differentiable 3D curve rendering pipeline which integrates DiffVG as a key component. Although 3Doodle achieves promising results for generating 3D sketches, its optimization requires supervision from multiview images, which makes it more like a reconstruction pipeline instead of a more flexible generation pipeline. Meanwhile, its performance is constrained by the inherent limitations of DiffVG’s original design. It relies on an approximate perspective projection to obtain the 2D sketch rendering and ignores the depth order of the curves, which may lead to color conflicts in the 3D sketch and limit its effectiveness on more complex tasks like colored sketch generation.
@@
In this paper, we present Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. Specifically, we represent the 3D sketch as a set of 3D rational Bézier curves and perform the perspective projection to obtain the 2D rational Bézier curves. Then, we propose a new differentiable rasterizer based on DiffVG to accurately render the projected 2D curves based on the depth order, so that more accurate occlusion relationships between curves can be modeled, especially for colored curves. Our framework supports end-to-end optimization of 3D curve primitives under flexible supervisions such as the distillation-based loss. By employing the recent Score Distillation Sampling algorithm to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the text or single image input. We conduct quantitative and qualitative comparisons between Diff3DS and related methods. The results demonstrate the superiority of our approach in generating view-consistent 3D sketches. Ablation studies further validate the effectiveness of the key components of our method. Additionally, analysis of the rasterizer shows its potential for generating colored sketches.
@@
We summarize our contributions as follows: (1) We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch from flexible inputs such as a single image or text. (2) We are the first to represent the 3D curve as 3D rational Bézier curve and design a depth-aware rasterizer that can enable precise and differentiable rendering of both black and colored 3D curves. (3) We conduct comprehensive experiments on novel text-to-3D and image-to-3D sketch generation tasks and the results demonstrate the superiority of our method.
@@ 2 Related Works
Existing research on 3D sketch generation mainly focus on interactive artistic sketch creation within 3D virtual systems (Wesche & Seidel, 2001; Grimm & Joshi, 2012; Bae et al., 2008), and immersive environments (Yu et al., 2021a; Xin et al., 2008; Jiang et al., 2021; Kim & Bae, 2016; Yu et al., 2021b; Arora & Singh, 2021). However, these works usually require users to possess professional artistic skills and 3D operation experience, making it less user-friendly for ordinary users. The recent 3Doodle (Choi et al., 2024) is able to generate expressive 3D sketches from multiview image observations, but it is more like multiview reconstruction instead of generation and the quality of the generated results highly depends on the number of observation viewpoints. By distilling prior knowledge from pre-trained 2D image generation models using methods like Score Distillation Sampling (Poole et al., 2023; Liu et al., 2023), we achieve user-friendly and view-consistent 3D sketch generation from flexible text or single image input.
@@
There have been several works for differentiable rendering of curves. DiffVG proposes a differentiable rasterizer for the creation of vector graphics represented by 2D curves and shapes. DRPG (Worchel & Alexa, 2023) investigates differentiable rendering of 3D parametric curves and surfaces, by piecewisely approximating the continuous parametric representations with a triangle mesh. 3Doodle (Choi et al., 2024) shares a similar differentiable rendering pipeline with us, including projecting the 3D Bézier curves to 2D Bézier curves, which are then rendered by the differentiable rasterizer from DiffVG. However, 3Doodle approximates the perspective projection as an orthographic projection to obtain 2D projected curves and ignores the depth ordering between the curves. Instead, we perform perspective projection to the 3D rational Bézier curves and designs a new DiffVG-based ratersizer which can maintain the depth ordering between the projected 2D rational Bézier curves.
@@
Motivated by the success of the text-to-image research, pioneering works (Poole et al., 2023; Wang et al., 2023a) introduce the Score Distillation Sampling (SDS) algorithm and leverage the pre-trained text-to-image models as prior knowledge to optimize 3D representations. Subsequent works combine SDS with various differentiable 3D representations to explore the capabilities of text-to-3D object generation, such as DMTet (Lin et al., 2023; Chen et al., 2023) and 3D Gaussians (Tang et al., 2024; Yi et al., 2024). Furthermore, Zero-1-to-3 (Liu et al., 2023) proposes an image-guided SDS algorithm that distills the 3D-consistent priors to optimize 3D representations given the input image, and its pipeline inspires the development of subsequent image-to-3D object methods (Liu et al., 2023; Qian et al., 2024; Wang & Shi, 2023; Wu et al., 2023; Sta, 2023). Distinguishing from prior endeavors, our work stands out by focusing on generating view-consistent 3D sketch by optimizing 3D parametric curve primitives.
@@ 3 Preliminary
DiffVG (Li et al., 2020) proposes a differentiable rasterizer for vector graphics that supports converting 2D Bézier curves to the raster image domain and back-propagating the gradient for optimization. Given the 2D curves parameters $\Theta$, the vector graphics scene is defined as $f(x,y,\Theta)$, and the raster image is defined as the 2D grid sampling over the space of $f(x,y,\Theta)$. To compute the color of a 2D location $(x,y)\in\mathbb{R}^{2}$, DiffVG utilizes the inside-outside test (Neh, ) to find the curves overlapped with the location initially. Subsequently, it sorts them according to a user-specified order and calculates the color using alpha blending (Porter & Duff, 1984). Due to the inside-outside test, the scene function $f$ is not differentiable with respect to curve parameters. Thus, it uses the anti-aliasing technique to make the pixel color differentiable. By prefiltering $f$ over a convolution kernel $k$ with support $A$, sampling $I(x,y)$ at pixels can yield an alias-free image:
@@
DiffVG proposes two approaches Monte Carlo Sampling and Analytical Prefiltering to evaluate the pixel integral, which does not have a closed-form solution in general. We focus on the first approach which performs better. The pixel integral can be discretized with Monte Carlo sampling (Ganacim et al., 2014):
@@ 4.1 Overview
We propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves. A 3D sketch $\tilde{\Theta}$ is defined as a collection of 3D parametric curve strokes that reside in 3D space. We particularly focus on the 3D rational Bézier curve and our derivation supports the 3D linear, quadratic, and cubic rational Bézier curves. The theoretical trainable parameters include the control point position, curve width, and rgb-alpha color. Follow the previous methods (Vinker et al., 2022; 2023; Choi et al., 2024), we represent each stroke by a 3D cubic Bézier curve and only optimize the control point position in our current implementation. Moreover, weights for rational Bézier curves are disregarded to simplify the computation. Fig. 3 illustrates the pipeline.
@@
In this section, we primarily explain how to project the 3D rational Bézier curves in camera coordinates onto the 2D camera focal plane, and detailed proof is provided in Appendix B. Given the control points $\tilde{P}_{i}=(\tilde{P}_{i}^{x},\tilde{P}_{i}^{y},\tilde{P}_{i}^{z})\in\mathbb{R}^{3}$, the associated non-negative weights $\tilde{w}_{i}$ and the Bernstein polynomials $B_{i,n}(t)$, for $0\leq t\leq 1$, define the 3D rational Bézier curve $\tilde{p}(t)$ of degree $n$ by:
@@
Due to rational Bézier curves are projective invariant, the curve $p(t)$ still is identical to the 2D rational Bézier curve defined by projected control points. Given projected 2D control points $P_{i}=(P_{i}^{x},P_{i}^{y})=(\frac{\tilde{P}_{i}^{x}}{\tilde{P}_{i}^{z}}f,\frac{\tilde{P}_{i}^{y}}{\tilde{P}_{i}^{z}}f)\in\mathbb{R}^{2}$ and adjusted weights $w_{i}=\tilde{w}_{i}\tilde{P}_{i}^{z}$, $p(t)$ can be written as following:
@@
This section primarily focuses on the differentiable rasterization of 2D curves $\Theta$ obtained through projection. To render the projected rational Bézier curves and maintain their occlusion relationship in 3D space, especially for colored curves, we customize a new differentiable rasterizer based on DiffVG (Li et al., 2020). Given a 2D pixel location $(x,y)$ and projected curves $\Theta$, we first employ a inside-outside test to identify all the overlapping curves with this pixel location. Then, we sort the overlapping curve points and calculate the color of this pixel location using alpha blending algorithm.
The inside-outside test identifies a curve as overlapping with a location if the closest distance between them is less than half of the curve width, and the closest point to the location are considered as the overlapping point . However, the identification is challenging. For the $n$th-degree rational curve $p(t)$, computing the $t$ that minimizes the distance $(p(t)-q)^{2}$ where $q=(x,y)$ equals to solve the roots of a polynomial with degree $23n-$ 2. For the cubic rational curve, a 7th order polynomial needs to be solved, which does not have a closed-form solution. Inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method (Press et al., 2007). The iterative solver obtains its initial guess from isolator polynomials (Sederberg & Chang., 1994) – the real roots of a 7th order polynomial can be isolated by the roots of two cubic auxiliary polynomials. Please refer to Appendix C for more details.
The rasterized result should faithfully maintains the occlusion in 3D space. In our task, different points along the same projected 2D curve may have different depths. Therefore, each point should be assigned with its own specific order for the color blending process. In the original DiffVG, each curve is initialized with a user-specific order, meaning all the points along the same curve share this uniform order. In contrast, our sort order is determined by the depth order among curve primitives. For each overlapping point, we compute its unprojected z-depth in the 3D space using the corresponding 2D projected curve control points, and employ it as the primary sort order. To mitigate the z-fighting issue  [^1] https://en.wikipedia.org/wiki/Z-fighting caused by floating-point errors, we also utilize a user-specified order as the secondary order. After obtaining the 2D curve scene function $f(x,y,\Theta)$, the rendered raster image $I$ will be calculated using the Monte Carlo Sampling strategy (Sec. 3) with Eq. 3. Fig. 2 shows an example. Our rasterizer faithfully renders the colored curves and maintains occlusions according to the depth order.
Figure 2: A rasterized result contains 3 quadratic rational Bézier curves and a line. All curves share the same control point positions in pixel space and the depths but with different weights. Our rasterizer faithfully renders the curves and maintains occlusions according to the depth order. (e.g., The upper half of each curve has a greater depth than the line, while the lower half has a lesser depth. This difference results in varying color blending outcomes at the overlapping regions).